### v1

@@ Abstract
Data-driven learning is rapidly evolving and places a new perspective on realizing state-space dynamical systems. However, dynamical systems derived from nonlinear ordinary differential equations (ODEs) suffer from limitations in computational efficiency. Thus, this paper stems from data-driven learning to advance states of dynamical systems utilizing a structured neural network (StNN). The proposed learning technique also seeks to identify an optimal, low-complexity operator to solve dynamical systems, the so-called Hankel operator, derived from time-delay measurements. Thus, we utilize the StNN based on the Hankel operator to solve dynamical systems as an alternative to existing data-driven techniques. We show that the proposed StNN reduces the number of parameters and computational complexity compared with the conventional neural networks and also with the classical data-driven techniques, such as Sparse Identification of Nonlinear Dynamics (SINDy) and Hankel Alternative view of Koopman (HAVOK), which is commonly known as delay-Dynamic Mode Decomposition(DMD) or Hankel-DMD. More specifically, we present numerical simulations to solve dynamical systems utilizing the StNN based on the Hankel operator beginning from the fundamental Lotka-Volterra model, where we compare the StNN with the LEarning Across Dynamical Systems (LEADS), and extend our analysis to highly nonlinear and chaotic Lorenz systems, comparing the StNN with conventional neural networks, SINDy, and HAVOK. Hence, we show that the proposed StNN paves the way for realizing state-space dynamical systems with a low-complexity learning algorithm, enabling prediction and understanding of future states.
keywords:
Dynamical systems, Neural Networks, Operator Learning, Data-driven Algorithms, Low-complexity Algorithms, Performance of Learning Algorithms, Nonlinear ODEs
{AMS}
34N05, 37L05, 65L20, 65Y20, 68T07, 68T07
1 Introduction
@@
The paper is organized as follows. We propose a simple structured operator called the Hankel operator and utilize that to determine time advance states in dynamical systems using efficient computations in Section 2. Next in Section Section 3 we propose to optimize the operator to realize time-advanced states of dynamical systems. In Section 4, we introduce a StNN designed to effectively represent states in dynamical systems. This section details the network’s architecture, highlighting its advantages over conventional NN. Furthermore, we provide an analysis of the flop counts and parameter efficiency of our proposed network. In the following Section, Section 5, we leverage the StNN to learn and update the states of the highly chaotic and extensively studied Lorenz system. This section will also showcase numerical results demonstrating the accuracy and precision of our proposed network. Additionally, we will conduct a long-term prediction analysis of our method, comparing its performance against conventional neural network predictions as well as those from SINDy and HAVOK, i.e., the delay-DM approach. Next, Section 5.1 presents results comparing StNN on an advanced approach for solving systems of ODEs across environments. Finally, we conclude the paper in Section 6.
@@
2 Preliminaries & Factorizations: Learning and Realizing States for Dynamical Systems
@@
2.1 Frequently Used Notations
@@
(1) ${\bf x}_{k}=\begin{bmatrix}x_{1}(t_{k})&x_{2}(t_{k})&\cdots&x_{n}(t_{k})\end{bmatrix}^{T},$
where T for the transpose, and $k=0,1,\cdots,n-1$. We utilize time-delays series of a state measurement $\{x(\tau_{k})\}_{k=0}^{n-1}$ to define a Hankel operator ${\bf H}$ s.t.
(2) ${\bf H}:=\begin{bmatrix}x(\tau_{0})&x(\tau_{1})&x(\tau_{2})&\cdots&x(\tau_{n-2})&x(\tau_{n-1})\\ x(\tau_{1})&x(\tau_{2})&x(\tau_{3})&\cdots&x(\tau_{n-1})&x(\tau_{0})\\ x(\tau_{2})&x(\tau_{3})&x(\tau_{4})&\cdots&x(\tau_{0})&x(\tau_{1})\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ x(\tau_{n-1})&x(\tau_{0})&x(\tau_{1})&\cdots&x(\tau_{n-3})&x(\tau_{n-2})\\ \end{bmatrix},$
where $\tau_{k}$’s are time-delay measurements. For a given vector ${\bf v}=[v_{0},v_{1},\cdots,v_{n-1}]^{T}\in R^{n}$, let us introduce an even-odd permutation matrix ${\bf P}_{n}$ ($n\geq 3$) by
${\bf P}_{n}\>\textbf{v}=\left\{\begin{array}[]{c}\left[v_{0},v_{2},\cdots,v_{n-2},v_{1},v_{3},\cdots,v_{n-1}\right]^{T}\>\textrm{even\>}n,\\ \left[v_{0},v_{2},\cdots,v_{N-1},v_{1},v_{3},\cdots,v_{n-2}\right]^{T}\>\textrm{odd\>}n.\end{array}\right.$
@@
2.2 Preliminaries: Dynamical Systems and Operator
@@
and it sees the states of the system at the $k^{\rm th}$ iteration as ${\bf x}_{k}\in\mathbb{R}^{n}$ having a non-linear function ${\bf F}$, which will usually denote iterations forward in time, so that ${\bf x}_{k}={\bf x}(k\Delta t)$. This is the situation, in which we could seek the solution of a dynamical system as a solution of a system of linear equations. Thus, to sum up, many problems in dynamical systems ultimately lead to a solution of systems of linear equations. On the other hand and due to the nonlinear nature of these dynamical systems, we propose a learning algorithm to train a neural network so that the network could learn an updated state from ${\bf x}_{k}$ (1) to ${\bf x}_{k+1}$ using a low-complexity ML operator. To learn an operator, we start with a vector-valued measurement function $g:M\rightarrow\mathbb{R}^{n}$ which are elements of an infinite-dimensional Hilbert space, i.e., Lebesgue square-integrable functions on a smooth manifold $M$. We define so-called a Hankel operator ${\bf H}$ (2) via time-delays series of a state measurement $\{x(\tau_{k})\}_{k=0}^{n-1}$ and extend it to states ${\bf x}_{k}$ by using time snapshots of spatiotemporal data that act on the measurement functions $g$ s.t. ${\bf H}g({\bf x})=(g\circ{\bf F})({\bf x})$, where $\circ$ is the composition operator and ${\bf F}$ is the non-linear function. More precisely for the discrete-time system with time-step $\Delta\tau$, the Hankel operator is defined as
(3) ${\bf H}g({\bf x}_{k})=g({\bf F}({\bf x}_{k}))=g({\bf x}_{k+1}).$
In other words, the proposed Hankel operator defines linear dynamical systems that advance the observations of the state ${\bf x}_{k}$, i.e., $g_{k}=g({\bf x}_{k})$ to the next time-advanced observations $g_{k+1}$ s.t.
$g({\bf x}_{k+1})={\bf H}g({\bf x}_{k}),$
for $k=0,1,\cdots,n-1$. We note that the Hankel operator is linear in the function spaces s.t. ${\bf H}(c_{1}g_{1}({\bf x})+c_{2}g_{2}({\bf x}))=c_{1}{\bf H}g_{1}({\bf x})+c_{2}{\bf H}g_{2}({\bf x})$, where $c_{1}$ and $c_{2}$ are constants and $g_{1}$ and $g_{2}$ are measurement functions.
Finally, we note that the Hankel operator for the dynamical systems evolved with continuous- time is defined as
$\frac{d}{dt}g({\bf x})={\bf H}g({\bf x}).$
@@
Thus, having a rich set of information based on the time-delayed operator ${\bf H}$ to predict future states of chaotic systems leads to better prediction than linear or nonlinear systems with trajectories trapped at fixed points or on periodic orbits [12]. On the other hand, instead of advancing linear or non-linear measurements of the states of a system, like in the DMD, we could measure time-delayed measurements using the Hankel operator ${\bf H}$ following the HAVOK [2, 10] and utilize that to obtain low-complexity algorithms to realize state measurements.
2.3 Factorize the Hankel Operator to Realize State Measurements
@@
Proposition 2.1.
Let ${\bf H}$ be the Hankel operator (2) determined via time-delays series of a state measurement $\{x(\tau_{k})\}_{k=0}^{n-1}$. Then, the Hankel operator can be utilized to advance the state ${\bf x}_{k}$ to the next time-advanced state ${\bf x}_{k+1}$ through the summation of the following low-rank matrices
(4) ${\bf x}_{k+1}$ $={\bf H}{\bf x}_{k}$
$={\bf J}^{T}[x_{1}(t_{k})\tilde{\bf x}+x_{2}(t_{k}){\bf Z}\tilde{\bf x}+x_{3}(t_{k}){\bf Z}^{2}\tilde{\bf x}+\cdots+x_{n}(t_{k}){\bf Z}^{n-1}\tilde{\bf x}],$
where ${\bf Z}$ is $r\times r$ upper shift matrix and
@@
Proof 2.2.
The operator ${\bf H}$ is a symmetric Hankel matrix determined by the first column(or row) of ${\bf H}$ s.t. $[x(\tau_{0}),x(\tau_{1}),\cdots,x(\tau_{n-1})]^{T}$, and when $Z$ is the lower shift matrix and when $\tilde{\bf x}_{r\times 1}$ is defined as above, we could write ${\bf H}=[\tilde{\bf x},Z\tilde{\bf x},\cdots,Z^{n-1}\tilde{\bf x}]$. Thus by taking ${\bf H}{\bf x}_{k}$ as a linear combination of the columns of ${\bf H}$ with scaling factors corresponding to the state ${\bf x}_{k}$ followed by scaling ${\bf J}^{T}$ and equation (3), we could compute ${\bf H}{\bf x}_{k}$ to realize ${\bf x}_{k+1}$ via (4).
Corollary 2.3.
Let the Hankel operator ${\bf H}$ be utilized to advance the observation of states from ${\bf x}_{k}$ to ${\bf x}_{k+1}$ using Proposition 2.1, then complexity in realizing time-advanced states cost $\mathcal{O}(n^{s})$, where $21<s<$ 2.
Proof 2.4.
Since ${\bf H}$ is the structured matrix determined by $\mathcal{O}(n)$ elements, we could compute ${\bf H}{\bf x}_{k}$ by utilizing the upper shift matrix ${\bf Z}_{r\times r}$ followed by the vector $\tilde{\bf x}_{r\times 1}$ in (4) to reduce the complexity in computing the conventional matrix-vector product of ${\bf H}{\bf x}_{k}$ from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n^{s})$, where $21<s<$ 2.
@@
Proposition 2.5.
Let ${\bf H}$ be the Hankel operator (2) determined via time-delays series of a state measurement $\{x(\tau_{k})\}_{k=0}^{n-1}$. Then, the time-advanced states ${\bf x}_{k+1}$ can be computed using the time-delayed states ${\bf x}_{k}$ followed by the decomposition of ${\bf H}$ s.t.
(5) ${\bf x}_{k+1}={\bf H}{\bf x}_{k}=\tilde{\bf I}_{n}\hat{\bf D}_{n}[{\bf J}^{T}]_{n\times r}{\bf C}_{r}[{\bf J}]_{r\times n}\hat{\bf D}_{n}\cdot{\bf x}_{k},$
where the circulant matrix ${\bf C}_{r}={\bf\mathfrak{F}}^{*}_{r}\breve{{\bf D}}_{r}{\bf\mathfrak{F}}_{r}$, and $\hat{D}_{n}=\left[diag(e^{-\frac{\pi k^{2}i}{n}})\right]_{k=0}^{n-1}$.
Proof 2.6.
When $n=2^{p}(p\geq 1)$, we could compute ${\bf H}{\bf x}_{k}$ by using the 2-FFTs as described in [37], and when $n\neq 2^{p}$, we could again compute ${\bf H}{\bf x}_{k}$ by using the 2-FFTs in [35].
Corollary 2.7.
Let the Hankel operator ${\bf H}$ be utilized to advance the observation of states from ${\bf x}_{k}$ to ${\bf x}_{k+1}$ using Proposition 2.5, then complexity in realizing time-advanced states is $\mathcal{O}(n\>\log n)$.
Proof 2.8.
As for any $n$, the product ${\bf H}{\bf x}_{k}$ could be computed using the 2-FFTs [37, 35], and hence the complexity in computing ${\bf H}{\bf x}_{k}$ to realize time advanced states cost $\mathcal{O}(n\>\log n)$.
3 A Best-fit Operator to Advance States
Following the previous section, we propose to obtain a best-fit operator for the Hankel operator ${\bf H}$–say $\widehat{\bf H}$ (w.l.o.g. let us assume that $m\geq n$), determined via time snapshots of spatiotemporal data. Furthermore, we propose to enhance learning by capturing the evolution of the nonlinear dynamical system using data-driven embedding based on an optimized operator.
3.1 Learn a Best-fit Operator
We first obtain a best-fit operator–say $\widehat{\bf H}$ determined via time-delays series of a state measurement $\{x(\tau_{k})\}_{k=0}^{n-1}$ to optimize the data-driven learning.
Proposition 3.1.
(6) ${\bf{X^{\prime}}}\approx\widehat{\bf H}{\bf X},\quad\widehat{\bf H}={\rm argmin}_{\bf H}\bigl{\{}\frac{1}{2}||{\bf{X^{\prime}}}-{\bf H}{\bf X}||^{2}_{F}+\alpha||{\bf H}||_{\eta}\bigr{\}},$
Proof 3.2.
(7) ${(\bf{X^{\prime}})^{T}}\approx{\bf X}^{T}\widehat{{\bf H}^{T}}\>\quad{\rm where},\>{\widehat{{\bf H}^{T}}}={\rm argmin}_{{\bf H}^{T}}\bigl{\{}\frac{1}{2}||({\bf{X^{\prime}}})^{T}-{\bf X}^{T}{\bf H}^{T}||^{2}_{F}+\alpha||\bf{\bf H}^{T}||_{\eta}\bigr{\}},$
(8) $\begin{split}&{\widehat{{\bf H}^{T}}}={\rm argmin}_{{\bf H}^{T}}\bigl{\{}||{\bf H}^{T}-\left({\bf H}^{T}_{k-1}-\frac{1}{t_{k}}{\bf X}({\bf X}^{T}{\bf H}^{T}_{k-1}-({\bf{X^{\prime}}})^{T})\right)||^{2}_{F}+\frac{2\alpha}{t_{k}}||\bf{\bf H}^{T}||_{\eta}\bigr{\}},\end{split}$
where ${\bf H}^{T}_{k-1}$ is the $k-1$ iterates for $H$ and $t_{k}$ is the stepsize. The minimization problem (8) can be solved by computing the singular value decomposition (SVD) of $\left({\bf H}^{T}_{k-1}-\frac{1}{t_{k}}{\bf X}({\bf X}^{T}{\bf H}^{T}_{k-1}-({\bf{X^{\prime}}})^{T})\right)$. Then the soft-thresholding operator can be applied on the singular values. By Theorem 2.1 in [15], the approximate solution for (8) has low-rank properties, which can be chosen as an approximate solution for the Hankel operator.
(9) ${\bf x}_{k+1}\approx\widehat{\bf H}{\bf x}_{k}+{\bf G}{\bf u}_{k},$
4 A Structured Neural Network (StNN) for Dynamical Systems

### v2

@@ Abstract
Data-driven learning is rapidly evolving and places a new perspective on realizing state-space dynamical systems. However, dynamical systems derived from nonlinear ordinary differential equations (ODEs) suffer from limitations in computational efficiency. To address this, we propose a structured neural network (StNN) that uses structured matrix theory and relies on a Hankel operator derived from time-delay measurements to solve dynamical systems. Specifically, the StNN identifies an optimal representation using the Hankel operator, providing a more computationally efficient alternative to existing data-driven approaches. We show that the proposed StNN places an optimal solution answering the demand for inference time, number of parameters, and computational complexity compared with the conventional neural networks and also with the classical data-driven techniques, such as Dynamic Mode Decomposition(DMD), Sparse Identification of Nonlinear Dynamics (SINDy), and Hankel Alternative View of Koopman (HAVOK), which is commonly known as delay-DMD or Hankel-DMD. Furthermore, we present numerical simulations to solve dynamical systems utilizing the StNN based on structured matrix theory followed by the Hankel operator beginning from the fundamental Lotka-Volterra model, where we compare the StNN with the LEarning Across Dynamical Systems (LEADS), and extend our analysis to highly nonlinear and chaotic Lorenz systems, comparing the StNN with conventional neural networks, DMD, SINDy, and HAVOK. Hence, we show that the proposed StNN paves the way for realizing state-space dynamical systems with a low-complexity learning algorithm, enabling prediction and understanding of future states.
Index Terms:
Dynamical systems, Structured Matrix Theory, Neural Networks, Operator Learning, Data-driven Algorithms, Low-complexity Algorithms, Performance of Algorithms, Nonlinear ODEs
I Introduction
@@
The paper is organized as follows. We propose a simple structured operator called the Hankel operator and utilize it to solve non-linear systems of ODEs using efficient computations in
@@
II Preliminaries & Factorizations: Learning and Realizing States for Dynamical Systems
@@
II-A Frequently Used Notations
@@
${\bf x}_{k}=\begin{bmatrix}x_{1}(t_{k})&x_{2}(t_{k})&\cdots&x_{n}(t_{k})\end{bmatrix}^{T},$ (1)
where T for the transpose, and $k=0,1,\cdots,n-1$. We utilize time-delays series of a state measurement $\{x(\tau_{k})\}_{k=0}^{n-1}$ to define a Hankel operator ${\bf H}\in\mathbb{R}^{n\times n}$ s.t.
${\bf H}:=\begin{bmatrix}x(\tau_{0})&x(\tau_{1})&x(\tau_{2})&\cdots&x(\tau_{n-2})&x(\tau_{n-1})\\ x(\tau_{1})&x(\tau_{2})&x(\tau_{3})&\cdots&x(\tau_{n-1})&x(\tau_{n-2})\\ x(\tau_{2})&x(\tau_{3})&x(\tau_{4})&\cdots&x(\tau_{n-2})&x(\tau_{n-3})\\ \vdots&\vdots&\vdots&\vdots&\mathinner{\mkern 2.0mu\raisebox{1.0pt}{.}\mkern 2.0mu\raisebox{4.0pt}{.}\mkern 2.0mu\raisebox{7.0pt}{.}\mkern 1.0mu}&\vdots\\ x(\tau_{n-2})&x(\tau_{n-3})&x(\tau_{n-4})&\cdots&x(\tau_{2})&x(\tau_{1})\\ x(\tau_{n-1})&x(\tau_{n-2})&x(\tau_{n-3})&\cdots&x(\tau_{1})&x(\tau_{0})\\ \end{bmatrix},$ (2)
where $\tau_{k}$’s are time-delay measurements. We note here that the matrix ${\bf H}$ (2) differs from HAVOK [7]. In HAVOK, the elements of Hankel matrix are defined based on an evaluation of states using the Koopman operator $\kappa$, specifically, the first column and row are defined as $[x(t_{1}),\kappa x(t_{1}),\cdots,\kappa^{q-1}x(t_{1})]$ and $[x(t_{1}),\kappa x(t_{1}),\cdots,\kappa^{p-1}x(t_{1})]^{T}$, respectively, where $p$ and $q$ are constants.
@@
II-B Preliminaries: Dynamical Systems and Operator
@@
and it sees the states of the system at the $k^{\rm th}$ iteration as ${\bf x}_{k}\in\mathbb{R}^{n}$ having a non-linear function ${\bf F}$, which will usually denote iterations forward in time, so that ${\bf x}_{k}={\bf x}(k\Delta t)$. This is the situation, in which we could seek the solution of a dynamical system as a solution of a system of linear equations. Thus, to sum up, many problems in dynamical systems ultimately lead to a solution of systems of linear equations. On the other hand and due to the nonlinear nature of these dynamical systems, we propose a learning algorithm to train a neural network so that the network could learn an updated state from ${\bf x}_{k}$ (1) to ${\bf x}_{k+1}$ using a low-complexity ML operator.
To learn an operator ${\bf H}$, we start with the discrete-time dynamical system ${\bf x}_{k+1}={\bf F}({\bf x}_{k})$ while defining the function evaluation as the matrix-vector computation via ${\bf F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ s.t. ${\bf F}({\bf x})={\bf H}{\bf x}$. Thus, the states of the system as it evolves in time can be defined via
${\bf x}_{k+1}={\bf F}({\bf x}_{k})={\bf H}{\bf x}_{k},$ (3)
We note that the Hankel operator is linear with respect to states $\bf x$ and $\tilde{\bf x}$ s.t. ${\bf H}(c_{1}{\bf x}+c_{2}\tilde{\bf x})=c_{1}{\bf H}{\bf x}+c_{2}{\bf H}\tilde{\bf x}$, where $c_{1}$ and $c_{2}$ are constants. Thus, to obtain the best-fit Hankel operator that best advances snapshot state measurements forward in time, we propose a best-fit operator as defined in the next section.
@@
Thus, having a rich set of information based on the time-delayed operator ${\bf H}$ to predict future states of chaotic systems leads to better prediction than linear or nonlinear systems with trajectories trapped at fixed points or on periodic orbits [7]. On the other hand, instead of advancing linear or non-linear measurements of the states of a system, like in the DMD, we could measure time-delayed measurements using the Hankel operator ${\bf H}$ following the HAVOK [32, 34] and utilize that to obtain low-complexity algorithms to realize state measurements as in next section.
II-C Learn a Best-fit Operator
We propose to obtain a best-fit operator for ${\bf H}$–say $\widehat{\bf H}$, determined via time snapshots of spatiotemporal data. Furthermore, we propose to enhance learning by capturing the evolution of the nonlinear dynamical system using data-driven embedding based on the best-fit operator.
Let us obtain the best-fit operator $\widehat{\bf H}$ determined via time-delays series of a state measurement $\{x(\tau_{k})\}_{k=0}^{n-1}$ to optimize the data-driven learning.
Proposition 1.
${\bf{X^{\prime}}}\approx\widehat{\bf H}{\bf X},\quad\widehat{\bf H}={\rm argmin}_{\bf H}\bigl{\{}\frac{1}{2}||{\bf{X^{\prime}}}-{\bf H}{\bf X}||^{2}_{F}+\alpha||{\bf H}||_{\eta}\bigr{\}},$ (4)
Proof.
$\begin{split}&{(\bf{X^{\prime}})^{T}}\approx{\bf X}^{T}\widehat{{\bf H}^{T}}\>\quad\\ &{\rm where},\\ &\>{\widehat{{\bf H}^{T}}}={\rm argmin}_{{\bf H}^{T}}\bigl{\{}\frac{1}{2}||({\bf{X^{\prime}}})^{T}-{\bf X}^{T}{\bf H}^{T}||^{2}_{F}+\alpha||\bf{\bf H}^{T}||_{\eta}\bigr{\}},\end{split}$ (5)
$\begin{split}&{\widehat{{\bf H}^{T}}}={\rm argmin}_{{\bf H}^{T}}\bigl{\{}||{\bf H}^{T}-\\ &\left({\bf H}^{T}_{k-1}-\frac{1}{t_{k}}{\bf X}({\bf X}^{T}{\bf H}^{T}_{k-1}-({\bf{X^{\prime}}})^{T})\right)||^{2}_{F}+\frac{2\alpha}{t_{k}}||\bf{\bf H}^{T}||_{\eta}\bigr{\}},\end{split}$ (6)
where ${\bf H}^{T}_{k-1}$ is the $k-1$ iterates for $H$ and $t_{k}$ is the stepsize. The minimization problem (6) can be solved by computing the singular value decomposition (SVD) of $\left({\bf H}^{T}_{k-1}-\frac{1}{t_{k}}{\bf X}({\bf X}^{T}{\bf H}^{T}_{k-1}-({\bf{X^{\prime}}})^{T})\right)$. Then the soft-thresholding operator can be applied on the singular values. By Theorem 2.1 in [37], the approximate solution for (6) has low-rank properties, which can be chosen as an approximate solution for the Hankel operator. ∎
${\bf x}_{k+1}\approx\widehat{\bf H}{\bf x}_{k}+{\bf G}{\bf u}_{k},$ (7)
II-D Factorize the Hankel Operator to Realize State Measurements
@@
Proposition 2.
Let ${\bf H}$ be the Hankel operator (2) determined via time-delays series of a state measurement $\{x(\tau_{k})\}_{k=0}^{n-1}$. Then, the Hankel operator can be calculated through the following low-rank matrices
${\bf H}={\bf H}_{l}+{\bf H}_{u}-x(\tau_{n-1})\tilde{\bf I}_{n},$
${\bf H}_{u}=[\tilde{\bf x},Z\tilde{\bf x},\cdots,Z^{n-1}\tilde{\bf x}],\>\>{\rm and}\>\>{\bf H}_{l}=\tilde{\bf I}_{n}[{\bf H}_{u}]^{T}\tilde{\bf I}_{n},$
where ${\bf Z}$ is $n\times n$ upper shift matrix and $\tilde{\bf x}=[x(\tau_{0}),x(\tau_{1}),x(\tau_{2}),\cdots,x(\tau_{n-1})]^{T}$
Proof.
The operator ${\bf H}$ is a per-symmetric Hankel matrix determined by the first column(or row) of ${\bf H}$ s.t. $[x(\tau_{0}),x(\tau_{1}),\cdots,x(\tau_{n-1})]^{T}$, and when $Z$ is the lower shift matrix and when $\tilde{\bf x}_{n\times 1}$ is defined as above, we could write ${\bf H}_{u}=[\tilde{\bf x},Z\tilde{\bf x},\cdots,Z^{n-1}\tilde{\bf x}]$ followed by the per-transpose to get ${\bf H}_{l}$ with the addition of $x(\tau_{n-1})\tilde{\bf I}_{n}$ to obtain ${\bf H}$. ∎
Corollary 1.
@@
Proof.
[…]
@@
Proposition 3.
Let ${\bf H}$ be the Hankel operator (2) determined via time-delays series of a state measurement $\{x(\tau_{k})\}_{k=0}^{n-1}$. Then, the operator can be realized using the following decomposition
${\bf H}=\tilde{\bf I}_{n}[{\bf J}^{T}]_{n\times r}{\bf C}_{r}[{\bf J}]_{r\times n},$ (8)
where the circulant matrix ${\bf C}_{r}={\bf\mathfrak{F}}^{*}_{r}\breve{{\bf D}}_{r}{\bf\mathfrak{F}}_{r}$.
Proof.
When $n=2^{p}(p\geq 1)$, we could compute ${\bf H}$ by using the 2-FFTs as described in [39], and when $n\neq 2^{p}$ we could pad with zeros to the nearest but the greatest power of 2 followed by the use of 2-FFTs in [39]. ∎
We note here that one could compute odd-length ${\bf H}$ using 2-FFTs as described in [43].
Corollary 2.
Let the Hankel operator ${\bf H}$ be utilized to advance the observation of states from ${\bf x}_{k}$ to ${\bf x}_{k+1}$ using Propositions 1 and 3, then complexity in realizing time-advanced states is $\mathcal{O}(n\>\log n)$.
Proof.
As for any $n$, the product ${\bf H}{\bf x}_{k}$ could be computed using the 2-FFTs [39, 43], and hence the complexity in computing ${\bf H}{\bf x}_{k}$ to realize time advanced states cost $\mathcal{O}(n\>\log n)$. ∎
III A Structured Neural Network (StNN) for Dynamical Systems